OpenStudy (anonymous):

how do you integrate x^2(rad(3x+4))

OpenStudy (anonymous):

\[\int\limits x^2/\sqrt{3x+4}\] ?

OpenStudy (anonymous):

your rad(3x+4) function makes no sense

OpenStudy (anonymous):

\[\int\limits x^{2}\sqrt{3x+4}\]

OpenStudy (anonymous):

Oh its multiplied.

OpenStudy (anonymous):

sorry for confusion

OpenStudy (anonymous):

Use substitution for the \[\sqrt{3x+4}\]

OpenStudy (anonymous):

What I did was: U=(3x+4)^(1/2) then solve for x. I got x=(1/3)(u^2-4). Find dx. dx=(1/3)(2u). Then plug it all in: \[\int\limits((1/3)u^2-4/3)(u)(2u/3)du=(2/9)\int\limits(u^2-4)(u^2)du.\] Straight forward from there. :)

OpenStudy (anonymous):

Miscopy. That should read:\[\int\limits((1/3)u^2-4/3)^2(u)(2u/3)du\]. That what you did myininaya?

myininaya (myininaya):

\[u=3x+4, du=3 dx\] \[x=\frac{u-4}{3}\] \[x^2=\frac{(u-4)^2}{9}\] \[\int\limits_{}^{} \frac{1}{3}*\frac{(u-4)^2}{9}*\sqrt{u} du\] \[\frac{1}{27}\int\limits_{}^{}(u^2-8u+16)*\sqrt{u} du\] \[\frac{1}{27}\int\limits_{}^{}(u^{\frac{5}{2}}-8u^{\frac{3}{2}}+16u) du\]

OpenStudy (anonymous):

That works too :P

OpenStudy (anonymous):

Mine was a tad more complicated looking haha

OpenStudy (anonymous):

where is the symbol for nice fractions like that myininaya?

myininaya (myininaya):

\[\frac{1}{27}*(\frac{2}{7}u^\frac{7}{2}-8*\frac{2}{5}u^\frac{5}{2}+16*\frac{1}{2}u^2)+C\]

myininaya (myininaya):

use frac{2}{3} will give you 2/3 in the equation editor

OpenStudy (anonymous):

nice thanks

OpenStudy (anonymous):

\[\frac{7}{x}\] I like it!!!

myininaya (myininaya):

we need to put that junk in terms of x

OpenStudy (anonymous):

yea lol make him do it I say

OpenStudy (anonymous):

Agreed^^

myininaya (myininaya):

i agree too :)